x+1/3y=-11
3/4x-1/2y=-15

Answer by emargo19(101)

Using the substitution method:-
x + 1/3y = -11 ----- equation 1
3/4x - 1/2y = -15 ----- equation 2
x = - 11 - 1/3y
3/4 ( -11 - y/3 ) - y/2 = - 15
- 33/4 - 3y/12 -y/2 = - 15
- 3y/12 - y/2 = -15 + 33/4
(-3y-6y)/12 = ( -60 + 33) / 4
-9y/12 = - 6.75
y = (-6.75 * 12)/-9
y = 9
Next lets substitute y=9
x + 1/3y = - 11 ----- equation 1
x + 3 = - 11
x = - 14